OpenStudy (anonymous):

i'm stuck on: integrate (cos(pi/x))/x^2

OpenStudy (anonymous):

-(sin(pi/x)/pi)+ some constant (C)

OpenStudy (anonymous):

i got the same answer. assume (pi/x) = u so du/dx = -pi/x^2 integral ((cos u)/x^2 ) . du x^2/-pi (-1/pi) integral (cos u) du = (-1/pi) sin u = (-1/pi) sin (pi/x)

OpenStudy (anonymous):

thank you this helps!

OpenStudy (anonymous):

you're welcome